# I Introduction

### v1

@@ VI.5 Rising towards a free surface: radial displacement field
VII Conclusions
@@
We performed falling-ball viscometry tests of hydrogel suspensions. We used hollow plastic spheres to observe the dynamics of a rising sphere as a function of time. The spheres rise with in cases with a free surface on the hydrogel suspension. Each run of the experiment has this feature. The ascent rate is found to depend exponentially on the buoyancy stress i.e. and therefore there is clear evidence for non-Newtonian behavior in the hydrogel suspension. We confirm the non-Newtonian behavior with temperature tests and flow field visualization, both of which reveal non-Stokesian dynamics. We contrast the linear rise dynamics with creep flow found earlier to confirm a dependence of the flow behavior on the boundary conditions. An intruder sinking towards a free boundary also shows linear displacement dynamics, and we find that when a freely floating lid is placed on the surface of the material, , with . The qualitative difference between the two types of motion is accompanied by clear evidence of flow at the free surface. The consistency of the results indicates that theoretical progress may be possible using continuum models in these flows that include boundary conditions as an effective way to tune the rheology of the bulk material.

### v2

@@ VI.5 Rising towards a free surface: radial displacement field
VII Discussion
Some intuition on the origin of the linear-to-sublinear transition of the intruder dynamics under the change of boundary conditions can be gained by considering a simple viscoelastic model to describe the hydrogel packing as a continuous fluid with an effective viscosity that sets the drag force on an intruder. Let be the effective buoyancy force experienced by the ping pong ball. In the overdamped limit in which we work, a simple force balance leads to the following differential equation
Even for a typical non-Newtonian rheology such as a Herschel-Bulkley fluid, which includes a yield stress, a constant motion perspective is to be expected, if the applied intruder stress is above the yield stress [13]. In this case, would be corrected for the existence of the yield stress and would be affected by the flow index.
In the free boundary case, thus, it seems reasonable to assume a constant viscosity , which leads to a constant velocity. We can then identify the exponential stress dependence of the hydrogel packing [22] as the mechanism to explain the thixotropic fluid response in the presence of a lid. Adding a lid at the top of the packing constrains or confines (at least partially) the hydrogel particles in the volume between the intruder and the lid itself, suppressing their motion and limiting their ability to escape from the aforementioned volume. Our hypothesis is that a modulation of the density in the packing, due to the combined effect of the lid confinement and of the intruder motion, is responsible for an increase in the total dissipation of intruder-induced hydrogel motion close to the rigid boundary. This, consequently, corresponds to a viscosity that grows with time, or, more importantly, that grows as the intruder approaches the rigid boundary, , allowing us to suggest a simple explanation of how the already known (but as yet unexplained) intrinsic rheological properties of the material could produce a time-dependent viscosity, and give rise to a sublinear motion of the intruder. We now develop this explanation step by step.
It was shown in previous studies [22] that the creep rate in hydrogel packings depends exponentially on the intruder stress, , and on the confinement stress acting on the hydrogel particles, . In the current study we confirm the exponential stress dependence in the linear case (cfr. section III) observing . Thus, we expect the dependence on the confinement stress to also hold: .
Our proposed mechanism is that this confinement stress, stored in the contact network of the hydrogel packing, increases as a result of the motion of the intruder towards a rigid boundary, . The intruder, effectively acts as a moving lid, pushing the hydrogel particles along on its trajectory towards the end boundary. The rigidity of the end boundary does not allow the particles to move radially as instead observed in the case of a free boundary (cfr section VI.5), so the internal stresses cannot relax. The hydrogels are, thus, constrained in a volume which cannot deform, the particles are compressed and the local density of the hydrogel packing increases (see Fig. 15 for a depiction of the mechanism). A perfect knowledge of the dependence of the excess confinement stress on the displacement of the ball, , would allow us to solve equation (3), where for small deviation of the confinement stress we could approximate the exponential stress dependence of the effective viscosity to the following
Assuming that , where is an exponent that depends on the end boundary condition and is a proportionality constant, we obtain which describes linear motion for and sublinear motion for . In particular, in the case of a linear dependence of the excess stress on the displacement, , the solution reads
which agrees with observations from previous studies of sinking intruders towards fully rigid boundaries. In the presence of a less constraining boundary condition, such as the floating lid considered in this study, there would be a sublinear dependence of the stress, and consequently of the viscosity on the travelled distance , and therefore the time history of the intruder will be intermediate between the linear free boundary case and the square-root rigid boundary case. Of course, our qualitative model does not substantiate the underlying mechanisms for the exponential stress dependence. We can also not experimentally access the packing fraction of the hydrogel packing, let alone gain clarity on the local variation of such density and their potential role in the viscosity enhancement. However, the quantitative potential of the model is clear.
VIII Conclusions
@@
We performed falling-ball viscometry tests of hydrogel suspensions. We used hollow plastic spheres to observe the dynamics of a rising sphere as a function of time. The spheres rise with in cases with a free surface on the hydrogel suspension. Each run of the experiment has this feature. The ascent rate is found to depend exponentially on the buoyancy stress i.e. and therefore there is clear evidence for non-Newtonian behavior in the hydrogel suspension. We confirm the non-Newtonian behavior with temperature tests and flow field visualization, both of which reveal non-Stokesian dynamics. We contrast the linear rise dynamics with creep flow found earlier to confirm a dependence of the flow behavior on the boundary conditions. An intruder sinking towards a free boundary also shows linear displacement dynamics, and we find that when a freely floating lid is placed on the surface of the material, , with . The qualitative difference between the two types of motion is accompanied by clear evidence of flow at the free surface. We hypothesize a quantitative connection between the change in boundary conditions, the exponential packing stress dependence and the time dependent flow behavior, stimulating theoretical progress while focusing future work on the yet unexplained but consistently observed exponential stress dependence of the flow behavior of hydrogel packings.